OpenStudy (anonymous):

the distributive property to find the value of 19 • 42 + 19 • 58 mentally

OpenStudy (anonymous):

You may notice that \(19\times 42+19\times 58=19\times (42+58)=19\times 100= 1900\). You calculated it mentally

OpenStudy (anonymous):

And Welcome to Openstudy